Find a polynomial function of least degree having only real​ coefficients, a leading coefficient of​ 1, and zeros of 5 and 3+i.
The polynomial function is ​f(x)=

Answer by ikleyn(52835)

.

This polynomial function is

    f(x) = (x-5)*(x-(3+i))*(x-(3-i)) = (x-5)*((x-3)-i)*((x-3)+i) = (x-5)*((x-3)^2+1) = (x-5)*(x^2-6x+10).    ANSWER

Solved.

-----------------

The lesson to learn


        You should include complex conjugate root with each complex root.

        The rest is simple arithmetic.